£20 is divided between Pip, Sara & Gordon so that Pip gets twice as much as Sara, and Sara gets three times as much as Gordon. How much does Pip get?
Answer by ikleyn(52781)

.

Let x = Gordon.

Then Sarah = 3x  and  Pip = 2*(3x) = 6x.


The total is


    x + 3x + 6x = 20,   or

    10x = 20  ===================>  x = 20/10 = 2.


ANSWER.  Pip get 6x = 6*2 = 12.

Solved.